OpenStudy (anonymous):

x over 6 = 7 over 2 you are suppose to cross multiply and I got, 2x=36. I then got the answer of 8 but it's wrong.. What did i do wrong?

OpenStudy (mathmale):

\[\frac{ x }{ 6 }=\frac{ 7 }{ 2 }\]

OpenStudy (mathmale):

Do cross multiply. What equation do you obtain from doing this?

OpenStudy (anonymous):

i am SO dumb. i would get 2x=42 !!! lol then the answer would be 21 !

OpenStudy (mathmale):

Right you are, and, by the way, you are NOT dumb!!

OpenStudy (anonymous):

thanks!

OpenStudy (mathmale):

Note that it helps quite a lot to write the problem out as shown: x 7 -- = --- 6 2 instead of x/6 = 7/2